# Reasoning with Confidence: Efficient Verification of LLM Reasoning Steps via Uncertainty Heads

### v1

@@ C.1 LLM Prompt
We use a domain-agnostic, format-enforcing prompt to elicit structured step-by-step reasoning, detailed in Figure 5.
@@
Figure 5: Prompt template used to elicit structured step-by-step reasoning from the model.
C.2 Annotation Prompts
For step-level annotation, we use the 2-stage prompts shown in Figure 6, while chain-level correctness annotations are obtained using the prompt in Figure 7.
@@ C.2 Annotation Prompts
You are given a problem, a ground-truth solution, and a step-by-step student solution. Your task is to analyze each step in the students solution to determine whether it is both logically correct and relevant.
@@
STUDENTS SOLUTION STEPS:
@@
Now, please evaluate whether the students steps are correct and logical.
@@
- A students step-by-step solution (as a Python list of string steps)
- An assessment of students solution
@@
STUDENTS SOLUTION STEPS:
@@
Figure 7: Prompt used for annotating chain-level correctness by evaluating the full reasoning trace.
Appendix D Training Details
D.1 Training Data Annotation Details
In our experiments, we employ two LLM judges for the annotation of UHead training data. For DeepSeek-R1, we follow the officially recommended inference setting, using a temperature of 0.6. Similarly, when using Qwen3-8B as training data annotator, we also follow the recommended inference hyperparameter, using a temperature of 0.7, a top_k of 20, and a top_p of 0.95. We access Qwen3-8B through vLLM local deployment, and access DeepSeek-R1 through the DeepSeek API. The annotation prompts are detailed in Figure 6. We do not set max tokens to restrict the length of reasoning chains.
D.2 Hyperparameter Details
For all experiments on Phi-4 and Qwen3-8B, we use the same set of hyperparameters. We use a learning rate of 5e-4, a batch size of 128, and a positive class weight of 3. For UHeads, we use one layer of transformer encoder with a hidden size of 512 and 16 attention heads. All training continues for 5 epochs. We use 10% of the training data and a 200-sample set of GSM8K, ScienceQA, and StrategyQA for validation and best checkpoint selection.
D.3 Computation Details
All experiments are carried out on cluster nodes with 4 GH200 GPUs and another cluster node with 2 H100 GPUs.
Appendix E Test Dataset Details
The mathematical domain includes MATH (high school and competition-style math problems; hendrycks2021measuringmathematicalproblemsolving), GSM8K (grade-school math word problems; cobbe2021trainingverifierssolvemath), and ProofNet (natural language proofs of undergraduate-level math problems; azerbayev2023proofnetautoformalizingformallyproving). Planning domain includes NaturalPlan (zheng2024naturalplanbenchmarkingllms) (spans three real-world tasks – trip planning, meeting planning, and calendar scheduling). The QA domain includes ScienceQA without multi-modal context (lu2022learn) (covers 26 science subjects from elementary to high school) and StrategyQA (geva2021strategyqa) – a general knowledge QA benchmark that requires implicit multi-step reasoning. Due to compute limitations, we evaluate on subsets of the test sets.
Table 10 presents the statistics of the test datasets used for the step-level benchmark. Importantly, we rely on DeepSeek-R1 for step-level evaluation, which is an expensive API-based LLM. We thus take a subset of test sets for this evaluation. While only having a few hundred questions, there is a substantial number of reasoning steps, which allows us to draw representative insights.
Table 10: Test dataset statistics used for step-level benchmark.
Dataset Model #Questions #Steps
Mean
#Steps
%Correct
@@ Appendix E Test Dataset Details
Answer Mean
Len (Tokens)
MATH Qwen3-8B 200 1203 6.2 85.5% 204
Phi-4 200 2296 11.5 89.2% 417
GSM8k Qwen3-8B 200 1056 5.1 94.0% 139
Phi-4 200 1286 6.4 95.7% 174
ProofNet Qwen3-8B 186 1211 6.5 81.9% 236
Phi-4 186 2227 12.0 91.1% 515
Trips Qwen3-8B 320 5492 17.2 46.9% 641
Phi-4 320 4486 14.0 44.87% 584
Meetings Qwen3-8B 200 2569 12.8 40.9% 739
Phi-4 200 3350 16.75 52.3% 584
Calendar Qwen3-8B 200 1616 8.1 52.5% 365
Phi-4 200 2172 10.86 66.9% 428
StrategyQA Qwen3-8B 500 2865 5.73 81.1% 121
Phi-4 500 3288 6.6 82.6% 182
ScienceQA Qwen3-8B 500 2621 5.24 88.7% 110
Phi-4 500 3415 6.8 91.9% 172
Appendix F Dataset Examples
Dataset Question Answer
- Step 2: Natalia sold half as many clips in May, which is $48\div 2=24$.
- Step 3: Add them: $48+24=72$.
<Answer>: 72
- Step 2: Then x=(r+x)rx=(r+x)-r is rational.
- Step 3: This contradicts $x$ being irrational.
- Step 4: Therefore $r+x$ is irrational.
<Answer>: $r+x$ is irrational
- Step 2: $\sqrt{50}\approx 7.071$.
- Step 3: Sum $\approx 12.548$.
- Step 4: Lies between 12 and 13.
<Answer>: 12 and 13
Trips You plan to visit 3 European cities for 14 days.
Stay in Istanbul for 5 days, Tallinn for 4 days, Zurich for 7 days.
From day 1-7, there is a show in Zurich.
Find a valid trip plan.
- Step 1: Start in Zurich for days 1-7 (show).
- Step 2: Fly Zurich Tallinn, stay days 8-11.
- Step 3: Fly Tallinn Istanbul, stay days 12-14.
<Answer>: Day 1-7: Zurich, Day 8-11: Tallinn, Day 12-14: Istanbul
Meetings You arrive at Financial District at 9:00AM.
Richard: Marina District 3:30-5:30PM (need 90 min).
Andrew: Alamo Square 5:00-10:00PM (need 120 min).
Travel: FD-Marina 15min, FD-Alamo 17min, Marina-Alamo 15min.
Plan a schedule to meet both.
- Step 1: Leave FD 2:45PM, arrive Marina 3:00PM.
- Step 2: Meet Richard 3:30-5:00PM.
- Step 3: Leave Marina 5:15PM, arrive Alamo 5:30PM.
- Step 4: Meet Andrew 5:30-7:30PM.
<Answer>: FD Marina Alamo schedule works
Laura busy: 11:30-12:30, 14:30-15:00, 16:00-16:30.
Paul busy: 9:30-10:00, 11:00-14:30, 15:00-17:00.
Paul prefers not after 9:30.
- Step 1: Laura free 9:00-11:30.
- Step 2: Paul free 9:00-9:30 and before 9:30 only.
- Step 3: Intersection is 9:00-9:30.
<Answer>: Meeting 9:00-9:30
- Step 2: Smartphones invented 21st century (iPhone 2007).
- Step 3: Impossible for him to have used one.
<Answer>: No, he did not
ScienceQA Which word is not like the others?
A. horse
B. goat
C. squirrel
D. leg
- Step 1: Horse, goat, squirrel are animals.
- Step 2: Leg is a body part.
- Step 3: Outlier is D.
<Answer>: D. leg
Table 11 presents examples of questions and LLM answers from the training datasets.
Appendix G Human Validation Details
We first randomly shuffle the DeepSeek-R1 annotated test set. The human annotation then starts from the first row of the shuffled dataset. Human annotators are not provided with DeepSeek-R1’s rationales (i.e., contents within the <think> tag) and must independently assess the correctness of each reasoning step, following the step-level correctness definition in Figure 6.
The annotation difficulty varies by tasks. ScienceQA is the easiest, which mostly requires common sense and usually takes 1 to 2 minutes to annotate one step. StrategyQA requires extensive fact-checking against Wikipedia, but since the dataset provides relevant facts for grading (available to both human and LLM annotators),annotation usually takes around 2 minutes per step. ProofNet and three planning datasets are the most demanding, often requiring more than 5 minutes per step due to the reasoning length and question complexity.
Appendix H The Usage of LLMs
In this paper, we mainly use LLMs as objects of study. We also use DeepSeek-R1 and Qwen3-8B as training data annotators, as detailed in § 3 and § C.1. We also use DeepSeek-R1 as a judge to grade answer correctness, as detailed in § 4. DeepSeek-R1’s accuracy in these tasks is manually validated through human annotation (see Table 1). In coding and writing, we use LLM assistants (e.g., ChatGPT) to identify grammar errors and debug. Such usage is under careful human supervision.

### v3

@@ F.1 Prompts for Reasoning
For LLMs in non-thinking mode, we use a domain-agnostic, format-enforcing prompt to elicit structured step-by-step reasoning, detailed in Fig 7. For thinking mode, we use a slightly different prompt specified in Fig 8.
@@
Figure 7: Prompt template for Phi-4 and Qwen3-8B in non-thinking mode. The prompt is designed to elicit structured step-by-step reasoning.
<|im_start|>user
Answer the following question. Enclose your answer in <answer></answer>.
<Question>: {q}<|im_end|>
<|im_start|>assistant
Figure 8: Prompt template for Qwen3-1.7B in native thinking mode.
F.2 Training Data Annotation Prompts
For step-level annotation, we use the 2-stage prompts shown in Fig 9, while chain-level correctness annotations are obtained using the prompt in Fig 10. Since the native thinking mode generates much longer CoT than in the non-thinking mode, we use a different annotation prompt. Specifically, rather than annotating all steps in one LLM call, we annotate one step per LLM call, using the question, ground truth, and previously generated steps as input. Fig 11 shows the annotation prompt template.
@@ F.2 Training Data Annotation Prompts
You are given a problem, a ground-truth solution, and a step-by-step student solution. Your task is to analyze each step in the student’s solution to determine whether it is both logically correct and relevant.
@@
STUDENT’S SOLUTION STEPS:
@@
Now, please evaluate whether the student’s steps are correct and logical.’
@@
- A student’s step-by-step solution (as a Python list of string steps)
- An assessment of student’s solution
@@
STUDENT’S SOLUTION STEPS:
@@
Figure 10: The prompt used for annotating chain-level correctness by evaluating the full reasoning trace.
You are given a problem, its ground-truth solution, and a student’s (incomplete) reasoning process so far.
Your task is to determine whether the **latest step** in the student’s reasoning is correct.
Instructions:
- A step is **wrong** if it contains explicit logical or computational errors, or if it contradicts any previous steps.
- Redundant, unnecessary, or non-informative steps are **not** considered wrong.
- If the latest step is correct, output **1**. If it is wrong, output **0**.
- Respond with the number (0 or 1) **only**, with no extra text or explanation.
PROBLEM:
{problem}
GROUND-TRUTH SOLUTION:
{answer}
STUDENT’S REASONING PROCESS SO FAR:
@@
THE LATEST STEP TO EVALUATE:
{latest_step}’
Figure 11: Prompt template for annotating correctness of reasoning steps in native thinking mode.
Appendix G The Usage of LLMs
In this paper, we mainly use LLMs as objects of study. We also use DeepSeek-R1 and Qwen3-8B as training data annotators, as detailed in § 3 and § F.1. We also use DeepSeek-R1 as a judge to grade answer correctness, as detailed in § 4. DeepSeek-R1’s accuracy in these tasks is manually validated through human annotation (see Table 13). In coding and writing, we use LLM assistants (e.g., ChatGPT) to identify grammar errors and debug. Such usage is under careful human supervision.